The current in a river was 3 mph. Jim could canoe downstream on this river for 92 miles in the same time it took him to canoe 68 miles upstream. How long did each trip take, and what was Jims rowing speed in the canoe?
Found 2 solutions by jorel555, ikleyn:
Answer by jorel555(1290)

Let n be Jim's rowing speed in still water. Then:
68/n-3=92/n+3
92(n-3)=68(n+3)
92n-276=68n+204
24n=480
n=20 mph in still water
68/17=4 hours rowing upstream
92/23=4 hours rowing downstream. ☺☺☺☺

Answer by ikleyn(52781)

.
For detailed solutions of similar problems see the lesson
    - Selected problems from the archive on the boat floating Upstream and Downstream
in this site.